OpenStudy (anonymous):

please help, which line is perpendicular to the line with the equation 2x+5y=20 A) y=--x B) y=--x+b C) y=-x+8 D) y=-x-29

OpenStudy (nincompoop):

perpendicular to a line means that the slope is going to be negative inverse so if you were given a line with a slope of 2/3, then the line perpendicular to it would be -3/2. another example is if the line given has a slope of -5/3, then the line perpendicular to it would have 3/5 slope

OpenStudy (nincompoop):

and why are you putting minus minus?

OpenStudy (nincompoop):

A) y=--x <--- are you sure that it is -- B) y=--x+b same with this one C) y=-x+8 D) y=-x-29

OpenStudy (anonymous):

yeah thats what it looks like on the paper, its a strait line

OpenStudy (nincompoop):

that is some typographical error. ask your teacher about it.

OpenStudy (anonymous):

shes not texting me back....

OpenStudy (nincompoop):

well it is sunday, what do you expect?

OpenStudy (anonymous):

ok so back to the question, so i understand that i would flip and add it but which one would i flip

OpenStudy (anonymous):

ment thay i would flip it like u said,,, ignore the add it part

OpenStudy (nincompoop):

what do you mean which one? the slope is what you obtain the reciprocal of, then make sure that it is negative of that slope \[\frac{ m }{ n }\rightarrow -\frac{ n }{ n }\] if your slope was originally negative, then the perpendicular line to it will be positive\[-\frac{ m }{ n } \rightarrow -\left(- \frac{ n }{ m } \right)=+\frac{ n }{ m }\]

OpenStudy (nincompoop):

the answer is not in the choices you were given

OpenStudy (anonymous):

ok thank u

OpenStudy (goformit100):

A warm Welcome to OpenStudy. I can help guide you through this useful site. You can ask your questions to me or you can message me. Please use the chat for off topic questions. Remember to give the person who helped you a medal by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it.